OpenStudy (anonymous):

How do you solve the equation 1,296^(2n)=216^(n+10)

OpenStudy (callisto):

\[1,296^{2n}=216^{n+10}\] \[(6^4)^{2n} = (6^3)^{n+10}\] \[6^{8n} = (6^{3n+30})\] Since both sides have the same base, you can rewrite it as \[8n = {3n+30}\] Can you solve it from here?

OpenStudy (anonymous):

no

OpenStudy (callisto):

Do you understand the above steps?

OpenStudy (anonymous):

Yes i just get stuck on the last step

OpenStudy (callisto):

Alright, -3n from both sides, what would you get?

OpenStudy (anonymous):

5 ?

OpenStudy (callisto):

Not really 8n - 3n = 3n+30 -3n What would you get?

OpenStudy (anonymous):

Im confused again

OpenStudy (callisto):

It's like solving a normal equation. Find the value of n...

OpenStudy (callisto):

8n - 3n = 3n+30 -3n 5n = 30 Divide both sides by 5, what would you get?